Question: One day, you find a wonderful money-making opportunity on the internet. Someone is willing to sell you a mineral called axinite for $9.00/kg. He furnished you with directions on how to extract aluminum from this mineral and also directs you to another company that is willing to pay you $50.00/kg for the aluminum you can provide. Assuming the procedure actually works as advertised, is this a good opportunity to get rich? You can assume that the formula for axinite is HCa3Al2BSi4O16.
1
Expert's answer
2020-08-12T08:35:39-0400

At first, we should find molar mass of axinite:

M(HCa3Al2BSi4O16)=1+3×40+2×27+11+4×28+16×16=1+120+54+11+112+256=554 (g/mol).

%Al=54/554×100%=9.75%.

So, from 1 kg (or 1000 g) of axinite we obtain 97.5 g of Al.

From a proportion we calculate how much we will get for them:

x=50×97.5/1000=4.875 $.

So, for 1 kg of axinite we pay 9$ and get only 4.875$ from it. It's obvious that this proposition is not good to get rich, because we spend more than we get.